# Optimal Spectral Transitions in High-Dimensional Multi-Index Models

### v1

@@ 3.3 Relation between the two spectral methods
As in the single-index setting, the tensors ${\underline{\bm{L}}}$ and ${\underline{\bm{T}}}$ are related by the following Proposition (proven in App. F).
@@
    Given an eigenpair $\gamma_{\underline{\bm{L}}}\geq 1,\bm{\Omega}$ of ${\underline{\bm{L}}}$, and defining ${\bm{W}}:={\bm{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}$, then:
    ${\underline{\bm{T}}}_{\gamma_{\underline{\bm{L}}}}\cdot{\bm{W}}\coloneqq\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\left(({\underline{\bm{G}}}+\gamma_{\underline{\bm{L}}}{\underline{\bm{I}}})^{-1}\cdot(\bm{\mathbf{X}}\bm{W})\right)={\bm{W}}.$ (42)
@@
    Given an eigenpair $\gamma_{\underline{\bm{T}}},{\bm{W}}$ of ${\underline{\bm{T}}}$ and defining
    $\bm{\Omega}:=({\underline{\bm{I}}}+{\underline{\bm{G}}})^{-1}\cdot({\bm{X}}{\bm{W}}),$ (43)
    then:
    ${\underline{\bm{L}}}\cdot\bm{\Omega}=\gamma_{\underline{\bm{T}}}\bm{\Omega}+(\gamma_{\underline{\bm{T}}}-1){\underline{\bm{G}}}\cdot\bm{\Omega}.$ (44)
Consequently, if there exists an eigenvector ${\bm{W}}$ of ${\underline{\bm{T}}}$ with eigenvalue $\gamma_{\underline{\bm{T}}}=1$, then $\bm{\Omega}$ in Eq. (43) is an eigenvector of ${\underline{\bm{L}}}$ with eigenvalue $\gamma_{{\underline{\bm{L}}}}=\gamma_{{\underline{\bm{T}}}}=1$.
@@ 4 Examples
Figure 4: Overlap $\|{\bm{M}}\|_{\mathrm{F}}^{2}/{\rm Tr}({\bm{Q}})$ as a function of the sample complexity $\alpha$. The dots represent numerical simulation results, computed for $n=5000$ (for the asymmetric method) or $d=5000$ (for the symmetric method) and averaging over $10$ instances. (Left) Link function $g(z_{1},z_{2})=z_{1}z_{2}$. Solid lines are obtained from state evolution predictions eq. (49,54). Dashed line at $\alpha_{c}\approx 0.59375$. (Right) Link function $g({\bm{z}})=p^{-1}\norm{{\bm{z}}}^{2}$, $p=4$. Solid lines are obtained from state evolution predictions eq. (47,54). Dashed line at $\alpha_{c}=2$.
@@
In this section we illustrate the framework introduced in Section 3 to predict the asymptotic performance of the spectral estimators (16,17) for specific examples of link functions, providing a comparison between our asymptotic analytical results and finite size numerical simulations for the overlap between the spectral estimators and the weights $\bm{W}_{\star}$, defined as $m\coloneqq\nicefrac{{\|{\bm{M}}\|_{\mathrm{F}}}}{{\sqrt{{\rm Tr}({\bm{Q}})}}}$, where ${\bm{M}}$ and ${\bm{Q}}$ are the overlap matrices defined in eq. (23) correspondent to the fixed points in Lemmas 3.3, 3.4, 3.8, 3.9. In Figure 4 we compare these theoretical predictions to numerical simulations at finite dimensions, respectively for the link functions $g(z_{1},z_{2})=z_{1}z_{2}$ and $g({\bm{z}})=p^{-1}\norm{{\bm{z}}}^{2}$. Additional numerical experiments are presented in Appendix D.
4.1 Asymmetric spectral method
@@
We provide closed-form expressions for the overlap parameter $m\coloneqq\nicefrac{{\|{\bm{M}}\|_{\mathrm{F}}}}{{\sqrt{{\rm Tr}({\bm{Q}})}}}$ of the spectral estimator $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{L}}}}$ (16), for a selection of examples of link functions. The details of the derivation are given in Appendix D.
  • $g(z\in\mathbb{R})$ (single-index model):
    $\alpha_{c}$ $=\left(\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\left(\text{Var}[z\big{|}{\mathrm{y}}]-1\right)^{2}\right]\right)^{-1},$ (45)
@@
    $m^{2}$ $=\Bigg{(}\frac{\alpha-\alpha_{c}}{\alpha+\alpha_{c}^{2}\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\left(\text{Var}[z\big{|}{\mathrm{y}}]-1\right)^{3}\right]}\Bigg{)}_{+}$ (46)
  • $g({\bm{z}})=p^{-1}||{\bm{z}}||^{2}$:
    $\alpha_{c}=\frac{p}{2},\quad m^{2}=\left(\frac{(2\alpha-p)}{2(\alpha+2)}\right)_{+}$ (47)
  • $g({\bm{z}})=\operatorname{sign}(z_{1}z_{2})$:
    $\alpha_{c}=\frac{\pi^{2}}{4},\quad m^{2}=\left(1-\frac{\pi^{2}}{4\alpha}\right)_{+}$ (48)
  • $g({\bm{z}})=\prod_{k=1}^{p}z_{k}$:
@@
    $\alpha_{c}$ $=\left(\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\lambda({\mathrm{y}})^{2}\right]\right)^{-1},$ (49)
    $m^{2}$ $=\left(\frac{\alpha-\alpha_{c}}{\alpha+\alpha_{c}^{2}\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\lambda({\mathrm{y}})^{3}\right]}\right)_{+},$ (50)
    where
@@
    $\lambda(y)=\begin{cases}\begin{array}[]{ll}|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}+{\mathrm{y}}-1,&p=2\\ \frac{2G^{p,0}_{0,p}\left(y^{2}2^{-p}\,\bigg{|}\,\begin{array}[]{c}0\\ \bm{e}_{p}\end{array}\right)}{G^{p,0}_{0,p}\left(y^{2}2^{-p}\,\bigg{|}\,\begin{array}[]{c}0\\ {\bm{0}}_{p}\end{array}\right)}-1,&p\geq 3\end{array}\end{cases}$ (51)
@@
4.2 Symmetric spectral method
We provide expressions for the overlap parameter $m\coloneqq\nicefrac{{\|{\bm{M}}\|_{\mathrm{F}}}}{{\sqrt{{\rm Tr}({\bm{Q}})}}}$ of the spectral estimator $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}$ (17), for a selection of examples of link functions. In all the following cases, the state evolution equations simplify, allowing to write the results as functionals of $\lambda:\mathbb{R}\to\mathbb{R}$, which is a specific to each problem:
  • $g(z\in\mathbb{R})$ (single-index model): $\lambda(y)=\text{Var}[z\big{|}y]-1$;
  • $g({\bm{z}})=p^{-1}||{\bm{z}}||^{2}$: $\lambda(y)=y-1$;
  • $g({\bm{z}})=\operatorname{sign}(z_{1}z_{2})$: $\lambda(y)=2\pi^{-1}y$;
  • $g({\bm{z}})=\prod_{k=1}^{p}z_{k}$: $\lambda(y)$ defined in eq. (51).
For all these example, the value $\alpha_{c}$ can be found in Section 4.1. For $\alpha>\alpha_{c}$, consider $a$ and $\gamma$ solutions of
$\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\frac{\lambda({\mathrm{y}})^{2}}{a(1+\lambda({\mathrm{y}}))-\lambda({\mathrm{y}})}\right]=\frac{1}{\alpha}$ (52)
$\gamma=1+\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\frac{\lambda({\mathrm{y}})}{a(1+\lambda({\mathrm{y}}))-\lambda({\mathrm{y}})}\right].$ (53)
@@ 4.2 Symmetric spectral method
which is strictly positive $\forall\alpha>\alpha_{c}$. Additional details on the derivation of this result can be found in Appendix E.
@@ 5 Conclusion and Perspectives
In this work, we tackled weak recovery in high-dimensional multi-index models via spectral methods, deriving two estimators inspired by a linearization of AMP. We showed they achieve the optimal reconstruction threshold, closing a key gap in prior approaches that required additional side information. Our analysis establishes that above the critical sample complexity, the leading eigenvectors of the proposed spectral operators align with the ground-truth subspace, echoing the BBP transition in random matrix theory.
This work advances our understanding of weak subspace recovery in multi-index models and provides a principled framework for designing optimal spectral estimators. It bridges ideas from random matrix theory, approximate message passing, and neural feature learning.
@@ Acknowledgements
The authors would like to thank Antoine Maillard, Benjamin Aubin and Lenka Zdeborova for early discussions on this problems in Santa Barbara. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), by the Swiss National Science Foundation under grant SNSF OperaGOST (grant number 200390) and by the French government, managed by the National Research Agency (ANR), under the France 2030 program with the reference "ANR-23-IACL-0008" and the Choose France - CNRS AI Rising Talents program.
@@ A.1 Linear GAMP
${\bm{g}}_{\rm out}^{t}(\bm{\omega},y)=\bm{G}^{t}(y)\bm{\omega},\quad\bm{f}_{in}^{t}({\bm{b}})={\bm{V}}^{t}{\bm{b}},$ (59)
@@
$\bm{\Omega}^{t}$ $=\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}^{t}-{\underline{\bm{G}}}^{t-1}\cdot\bm{\Omega}^{t-1}{\bm{V}}^{T}_{t},$ (60)
$\bm{\mathbf{\hat{W}}}^{t+1}$ $=\left(\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}^{t}-\bm{\mathbf{\hat{W}}}^{t}{\bm{A}}_{t}^{T}\right){\bm{V}}_{t+1}^{T},$ (61)
where $G_{ikjh}=\delta_{ij}[\bm{G}(y_{i})]_{kh}$ and
${\bm{A}}=d^{-1}\sum_{i\in\llbracket n\rrbracket}\bm{G}(y_{i})\xrightarrow{n,d\to\infty}\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}[\bm{G}({\mathrm{y}})].$ (62)
@@
$\bm{G}(y)=\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}},\quad{\bm{V}}={\bm{I}}$ (63)
@@
${\bm{A}}=\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}[{\bm{G}}({\mathrm{y}})]=\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}\big{|}{\mathrm{y}}]-\alpha{\bm{I}}=\alpha\mathbb{E}_{\bm{\mathbf{z}}\sim\mathsf{N}({\bm{0}},{\bm{I}})}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}]-\alpha{\bm{I}}={\bm{0}}.$ (64)

### v2

@@ 3.3 Relation between the two spectral methods
As in the single-index setting, $\bm{L}$ and $\bm{T}$ are related by the following Proposition (proven in App. G).
@@
Define $\hat{{\bm{G}}}\in\mathbb{R}^{np\times np}$ such that $\hat{G}_{(i\mu),(j\nu)}:=\delta_{ij}G_{\mu\nu}({\mathrm{y}}_{i})$.
@@
    Given an eigenpair $\gamma_{\bm{L}}\geq 1,\bm{\omega}\in\mathbb{R}^{np}$ of $\bm{L}$, if ${\bm{w}}:={\rm vec}\left(\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{\bm{G}}\bm{\omega}\right)\right)$, then
    ${\bm{T}}_{\gamma_{\bm{L}}}{\bm{w}}={\bm{w}},$ (31)
    with $\bm{T}_{\gamma}\in\mathbb{R}^{dp\times dp}$ defined as
    $[\bm{T}_{\gamma}]_{(k\mu),(h\nu)}:=\sum_{i\in\llbracket n\rrbracket}{\mathrm{X}}_{ik}{\mathrm{X}}_{ih}\left[{\bm{G}}({\mathrm{y}}_{i})\left({\bm{G}}({\mathrm{y}}_{i})+\gamma{\bm{I}}\right)^{-1}\right]_{\mu\nu}\quad k,h\in\llbracket d\rrbracket,\,\mu,\nu\in\llbracket p\rrbracket\,,$ (32)
@@
    Given an eigenpair $\gamma_{\bm{T}},{\bm{w}}\in\mathbb{R}^{dp}$ of $\bm{T}$ and defining $\bm{\omega}:=({\bm{I}}_{np}+\hat{\bm{G}})^{-1}{\rm vec}\left({\bm{X}}{\rm mat}({\bm{w}})\right)$, then
    $\bm{L}\bm{\omega}=\gamma_{\bm{T}}\bm{\omega}+(\gamma_{\bm{T}}-1)\hat{\bm{G}}\bm{\omega}.$ (33)
Consequently, if there exists an eigenvector ${\bm{w}}$ of $\bm{T}$ with eigenvalue $\gamma_{\bm{T}}=1$, then $\bm{\omega}=({\bm{I}}_{np}+\hat{\bm{G}})^{-1}{\rm vec}\left({\bm{X}}{\rm mat}({\bm{w}})\right)$ is an eigenvector of $\bm{L}$ with eigenvalue $\gamma_{\bm{L}}=\gamma_{\bm{T}}=1$.
@@ 4 Examples
Figure 4: Overlap $\|{\bm{M}}\|_{\mathrm{F}}^{2}/{\rm Tr}({\bm{Q}})$ as a function of the sample complexity $\alpha$. The dots represent numerical simulation results, computed for $n=5000$ (for the asymmetric method) or $d=5000$ (for the symmetric method) and averaging over $10$ instances. (Left) Link function $g(z_{1},z_{2})=z_{1}z_{2}$. Solid lines are obtained from state evolution predictions (Section 4). Dashed line at $\alpha_{c}\approx 0.59375$. (Right) Link function $g({\bm{z}})=p^{-1}\norm{{\bm{z}}}^{2}$, $p=4$. Solid lines are obtained from state evolution predictions (Section 4). Dashed line at $\alpha_{c}=2$.
@@
In this section we illustrate the framework introduced in Section 3 to predict the asymptotic performance of the spectral estimators (10,12) for specific examples of link functions, providing a comparison between our asymptotic analytical results and finite size numerical simulations for the overlap between the spectral estimators and the weights $\bm{W}_{\star}$, defined as $m\coloneqq\nicefrac{{\|{\bm{M}}\|_{\mathrm{F}}}}{{\sqrt{{\rm Tr}({\bm{Q}})}}}$, where ${\bm{M}}$ and ${\bm{Q}}$ are the overlap matrices defined in eq. (16) correspondent to the fixed points in Theorems 3.3, 3.4, 3.8, 3.9. In Figure 4 we compare these theoretical predictions to numerical simulations at finite dimensions, respectively for the link functions $g(z_{1},z_{2})=z_{1}z_{2}$ and $g({\bm{z}})=p^{-1}\norm{{\bm{z}}}^{2}$. Additional numerical experiments are presented in Appendix E.
Asymmetric spectral method.
@@
We provide closed-form expressions for the overlap parameter $m\coloneqq\nicefrac{{\|{\bm{M}}\|_{\mathrm{F}}}}{{\sqrt{{\rm Tr}({\bm{Q}})}}}$ of the spectral estimator $\bm{\mathbf{\hat{W}}}_{\bm{L}}$ (10), for a selection of examples of link functions. The details of the derivation are given in Appendix E.
@@
  • $g(z\in\mathbb{R})$ (single-index): $\alpha_{c}=\left(\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\left(\mathbb{E}[z^{2}-1|{\mathrm{y}}]\right)^{2}\right]\right)^{-1}$, $m^{2}=\Bigg{(}\frac{\alpha-\alpha_{c}}{\alpha+\alpha_{c}^{2}\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\left(\mathbb{E}[z^{2}-1|{\mathrm{y}}]\right)^{3}\right]}\Bigg{)}_{+}$
  • $g({\bm{z}})=p^{-1}||{\bm{z}}||^{2}$: $\alpha_{c}=p/2,\quad m^{2}=\left({(\alpha-\nicefrac{{p}}{{2}})(\alpha+2)^{-1}}\right)_{+}$
  • $g({\bm{z}})=\operatorname{sign}(z_{1}z_{2})$: $\alpha_{c}=\pi^{2}/4,\quad m^{2}=\left(1-\frac{\pi^{2}}{4\alpha}\right)_{+}$
@@
  • $g({\bm{z}})=\prod_{k=1}^{p}z_{k}$: $\alpha_{c}=\left(\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\lambda({\mathrm{y}})^{2}\right]\right)^{-1},\quad m^{2}=\left({(\alpha-\alpha_{c})\left(\alpha+\alpha_{c}^{2}\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\lambda({\mathrm{y}})^{3}\right]\right)^{-1}}\right)_{+},$
    where we used
@@
    $\lambda(y)=|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}+{\mathrm{y}}-1,\;\;(p=2)\quad\text{and}\quad\lambda(y)=\frac{2G^{p,0}_{0,p}\left(y^{2}2^{-p}\,\bigg{|}\,\begin{array}[]{c}0\\ \bm{e}_{p}\end{array}\right)}{G^{p,0}_{0,p}\left(y^{2}2^{-p}\,\bigg{|}\,\begin{array}[]{c}0\\ {\bm{0}}_{p}\end{array}\right)}-1,\;\;(p\geq 3)$ (34)
@@
Symmetric spectral method.
We provide expressions for the overlap parameter $m\coloneqq\nicefrac{{\|{\bm{M}}\|_{\mathrm{F}}}}{{\sqrt{{\rm Tr}({\bm{Q}})}}}$ of the spectral estimator $\bm{\mathbf{\hat{W}}}_{\bm{T}}$ (12), for a selection of examples of link functions. In all the following cases, the state evolution equations simplify, allowing to write the results as functionals of $\lambda:\mathbb{R}\to\mathbb{R}$, specific to each problem:
  • $g(z\in\mathbb{R})$: $\lambda(y)=\text{Var}[z\mid y]-1$ (single-index model);
  • $g(\mathbf{z})=\operatorname{sign}(z_{1}z_{2})$: $\lambda(y)=\tfrac{2}{\pi}\,y$
  • $g(\mathbf{z})=p^{-1}\|\mathbf{z}\|^{2}$: $\lambda(y)=y-1$;
  • $g(\mathbf{z})=\prod_{k=1}^{p}z_{k}$: $\lambda(y)$ defined in eq. (34).
For all these examples, the value $\alpha_{c}$ has been reported in the previous paragraph. For $\alpha>\alpha_{c}$, consider $a$ and $\gamma$ solutions of
$\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\frac{\lambda({\mathrm{y}})^{2}}{a(1+\lambda({\mathrm{y}}))-\lambda({\mathrm{y}})}\right]=\frac{1}{\alpha},\qquad\gamma=1+\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\frac{\lambda({\mathrm{y}})}{a(1+\lambda({\mathrm{y}}))-\lambda({\mathrm{y}})}\right].$ (35)
@@
which is strictly positive $\forall\alpha>\alpha_{c}$. In all the example the principal eigenvector is $\lambda_{s}=1$. Additional details can be found in Appendix F.
@@ 5 Conclusion and Perspectives
In this work, we tackled weak recovery in high-dimensional multi-index models via spectral methods, deriving two estimators inspired by a linearization of AMP. We showed they achieve the optimal reconstruction threshold, closing a key gap in prior approaches that required additional side information. Our analysis establishes that above the critical sample complexity, the leading eigenvectors of the proposed spectral operators align with the ground-truth subspace, echoing the BBP transition in random matrix theory. This work advances our understanding of weak subspace recovery in multi-index models and provides a principled framework for designing optimal spectral estimators. It bridges ideas from random matrix theory, approximate message passing, and neural feature learning.
@@ Acknowledgements
The authors would like to thank Antoine Maillard, Benjamin Aubin and Lenka Zdeborova for early discussions on this problems in Santa Barbara. We would also like to thank Filip Kovačević, Yihan Zhang and Marco Mondelli for the discussions on the relationship with their work following the first version of this manuscript. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), by the Swiss National Science Foundation under grant SNSF OperaGOST (grant number 200390) and by the French government, managed by the National Research Agency (ANR), under the France 2030 program with the reference "ANR-23-IACL-0008" and the Choose France - CNRS AI Rising Talents program.
@@ A.1 Linear GAMP
${\bm{g}}_{\rm out}^{t}(\bm{\omega},y)=\bm{G}_{\rm out}^{t}(y)\bm{\omega},\quad\bm{f}_{in}^{t}({\bm{b}})={\bm{V}}^{t}{\bm{b}},$ (41)
@@
$\bm{\Omega}^{t}$ $=\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}^{t}-{\bm{\Gamma}}_{\rm out}^{t-1}{\bm{V}}^{T}_{t},$ (42)
$[{\bm{\Gamma}}_{\rm out}^{t}]_{i\mu}$ $=\sum_{\nu\in\llbracket p\rrbracket}[{\bm{G}}_{\rm out}^{t}]_{\mu\nu}({\mathrm{y}}_{i})\Omega^{t}_{i\nu},\quad i\in\llbracket n\rrbracket,\,\mu\in\llbracket p\rrbracket,$ (43)
$\bm{\mathbf{\hat{W}}}^{t+1}$ $=\left(\bm{\mathbf{X}}^{T}{\bm{\Gamma}}_{\rm out}^{t}-\bm{\mathbf{\hat{W}}}^{t}{\bm{A}}_{t}^{T}\right){\bm{V}}_{t+1}^{T},$ (44)
where
${\bm{A}}=d^{-1}\sum_{i\in\llbracket n\rrbracket}\bm{G}_{\rm out}({\mathrm{y}}_{i})\xrightarrow{n,d\to\infty}\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}[\bm{G}_{\rm out}({\mathrm{y}})].$ (45)
@@
$\bm{G}_{\rm out}(y)=\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}=y],\quad{\bm{V}}={\bm{I}}$ (46)
@@
${\bm{A}}=\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}[{\bm{G}}_{\rm out}({\mathrm{y}})]=\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}\big{|}{\mathrm{y}}]-\alpha{\bm{I}}=\alpha\mathbb{E}_{\bm{\mathbf{z}}\sim\mathsf{N}({\bm{0}},{\bm{I}})}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}]-\alpha{\bm{I}}={\bm{0}}.$ (47)